A LED company claims that the average life of the LED light bulbs it

manufactures is 1, 500 hours with a standard deviation of 500 hours. If a random sample of 40 bulbs is chosen, what is the probability that the sample mean will be:

a. greater than 1, 400 hours?

b. less than 1, 400 hours?


1
Expert's answer
2021-05-31T10:41:14-0400

"z=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{1400-1500}{500/\sqrt{40}}=-1.26"

a)

"(\overline{x}>1400)=1-P(z<-1.26)=1-0.10565=0.89435"


b)

"(\overline{x}<1400)=P(z<-1.26)=0.10565"